# LoRA-Loop: Closing the Synthetic Replay Cycle for Continual VLM Learning

### v1

@@ 1 Introduction
Zheng et al.’s ZSCL [68] first tackled CL for VLMs under multi-domain settings by distilling on large, diverse reference data (e.g., ImageNet [47] or CIFAR [29]), aligning post-finetuning activations with their original values to prevent drift. They further proposed the more challenging and realistic Multi-domain Task Incremental Learning (MTIL) benchmark to evaluate performance across varied domains. More recently, GIFT [56] leverages a frozen Stable Diffusion model [46] to generate synthetic “reference” images on the fly, instead of storing a large real dataset. This approach cuts storage costs and enables concept rehearsal without any generator finetuning.
@@ 3.2.1 Synthetic Replay Sample Filtering for Distillation
and sort $\mathrm{conf}_{j}$ to retain the top‐$k$ pairs for each class to form $S$. This ensures $S$ contains the samples best aligned with the domain of the previous task data, improving distillation efficiency while controlling memory and compute.
@@ 4.1 Experiment Settings
Implementation Details. We build on prior continual VLM learning work [68, 56] using CLIP with a ViT‐B/16 backbone [8]. Each task is finetuned for $1{,}000$ iterations with a batch size of $64$. For synthetic replay, at each task we draw $M_{\rm pre}=8$ candidates per class using Stable Diffusion v1.5 [46] (classifier‐free guidance scale $7.5$, $50$ denoising steps) and retain the top-$k=1$ images per class based on CLIP cosine similarity. After updating the VLM to $f^{i}$, we score all training examples and select $l=2$ samples per class to form the LoRA training set. We then finetune a rank-$r=4$ LoRA adapter on Stable Diffusion for $100$ epochs using AdamW [37] with a learning rate of $1\times 10^{-4}$, $(\beta_{1},\beta_{2})=(0.9,0.999)$, and weight decay $1\times 10^{-2}$.
@@ 4.2.2 Ablation Studies
To assess hyperparameter sensitivity, we test five key settings and report results in Tab. 4: LoRA adapter rank $r$, number of per-class tuning examples $l$ at $r=4$ and $r=16$, LoRA training set selection policy, sample-filtering policy, and pre-filter sampling budget $M_{pre}$. As shown in Tab. 4(a), performance peaks at $r=4$, and increasing $r$ beyond causes a steady drop in all metrics, which we attribute to overly aggressive learning that degrades generation quality and thus weakens synthetic replay. With a moderate rank of $r=4$ (Tab. 4(b)), just two examples per class suffice to reach peak alignment, whereas larger $l$ values yield marginal declines, likely because finite capacity cannot absorb too many samples. In the case of $r=16$ (Tab. 4(c)), adding more tuning data neither improves stability nor boosts accuracy, suggesting that aggressive adaptation with limited prompt diversity can destabilize generation. Turning to selection policies, our Top & Bottom training set selection scheme (Tab. 4(d)) outperforms both random sampling and top-only confidence by inclusively covering both prototypical and edge cases during LoRA finetuning. In the sample filtering stage (Tab. 4(e)), retaining only the highest-confidence generations delivers the best overall performance, while including mid- or low-confidence outputs noticeably degrades results. Finally, increasing the pre-filter budget (Tab. 4(f)) steadily boosts performance up to $8$ samples per class, reflecting a more thorough search, but plateaus beyond that point. Overall, our experiments dissect each design choice, showing that LoRA finetuning and sample filtering jointly bridge domain and semantic gaps while stabilizing generation, achieving strong knowledge retention and zero-shot generalization across varied hyperparameters.

### v2

@@ 1 Introduction
Zheng et al.’s ZSCL [68] first tackled CL for VLMs under multi-domain settings by distilling on large, diverse reference data (e.g., ImageNet [47] or CIFAR [29]), aligning post-finetuning activations with their original values to prevent drift. They further proposed the more challenging and realistic Multi-domain Task Incremental Learning (MTIL) benchmark to evaluate performance across varied domains. More recently, GIFT [56] leverages a frozen Stable Diffusion (SD) model [46] to generate synthetic “reference” images on the fly, instead of storing a large real dataset. This approach cuts storage costs and enables concept rehearsal without any generator finetuning.
@@ 3.2.1 Synthetic Replay Sample Filtering for Distillation
[…]
@@ 4.1 Experiment Settings
[…]
@@ 4.2.2 Ablation Studies
To assess hyperparameter sensitivity, we test five key settings and report results in Tab. 4: LoRA adapter rank $r$, number of per-class tuning examples $l$ at $r=4$ and $r=16$, LoRA training set selection policy, sample-filtering policy, and pre-filter sampling budget $M_{pre}$. As shown in Tab. 4(a), performance peaks at $r=4$, and increasing $r$ beyond causes a steady drop in all metrics, which we attribute to overly aggressive learning that degrades generation quality and thus weakens synthetic replay. With a moderate rank of $r=4$ (Tab. 4(b)), just two examples per class suffice to reach peak alignment, whereas larger $l$ values yield marginal declines, likely because finite capacity cannot absorb too many samples. In the case of $r=16$ (Tab. 4(c)), adding more tuning data neither improves stability nor boosts accuracy, suggesting that aggressive adaptation with limited prompt diversity can destabilize generation. Turning to selection policies, our Top & Bottom training set selection scheme (Tab. 4(d)) outperforms both random sampling and top-only confidence by inclusively covering both prototypical and edge cases during LoRA finetuning. In the sample filtering stage (Tab. 4(e)), retaining only the highest-confidence generations delivers the best overall performance, while including mid- or low-confidence outputs noticeably degrades results. Finally, increasing the pre-filter budget (Tab. 4(f)) steadily boosts performance up to 888 samples per class, reflecting a more thorough search, but plateaus beyond that point. Overall, our experiments dissect each design choice, showing that LoRA finetuning and sample filtering jointly bridge domain and semantic gaps while stabilizing generation, achieving strong knowledge retention and zero-shot generalization across varied hyperparameters.